### v1

# sample.bib

Abstract.
A well-known result in random matrix theory, proven by Kahn, Komlós and Szemerédi in 1995, states that a square random matrix with i.i.d. uniform {±$\{\pm 1\}$ entries is invertible with probability $1-\exp(-\Omega(n))$. As a natural generalization of the model and motivated by the Anderson localization-delocalization transition in mathematical physics, we consider the invertibility of a class of random band matrices with independent entries where the bandwidth $d_{n}$ scales like $n^{\alpha}$, for some $\alpha\in(0,1)$. The band matrix model we consider is sufficiently general and covers existing models such as the block band matrix and periodic band matrix, allowing great flexibility in the variance profile. As the bandwidth is sublinear in the dimension, estimating the invertibility and least singular values of these matrices is a well-known open problem. We make a progress towards the invertibility problem by showing that, when $\alpha>\frac{2}{3}$ and when the random variables are i.i.d. uniformly distributed on {±$\{\pm 1\}$, then the band matrix is invertible with probability $1-\exp(-\Omega(n^{\alpha/2}))$. Previously, even invertibility with probability $1-o(1)$ was not known for these band matrix models except in the very special case of block band matrices.
1. Introduction
In this paper, we investigate the invertibility for a wide class of combinatorial random matrices with an inhomogeneous variance profile. Before stating the main result on band matrices, we begin with a historical account on the invertibility of $n\times n$ random matrices with i.i.d. entries, which is much better understood in the past thirty years. Let $M_{n}$ be an $n\times n$ matrix with each entry independently and uniformly distributed on $\{-1,1\}$, and denote by $P_{n}:=\mathbb{P}(\det M_{n}=0)$. The first proof of asymptotic non-singularity, that $\lim_{n\to\infty}P_{n}=0$, was given by Komlós [komlos1967determinant]. Then Kahn, Komlós and Szemerédi [kahn1995probability] provided the first exponential bound on $P_{n}$: $P_{n}\leq c^{n}$ for $n$ sufficiently large, and where they can take $c=0.999$. Tao and Vu [tao2007singularity] improved the value of $c$ to be $c=3/4+o(1)$, and Bourgain, Vu, Wood [bourgain2010singularity] further improved to $c=\sqrt{2}/2+o(1)$. Finally, Tikhomirov [tikhomirov2020singularity] obtained the optimal value $c=\frac{1}{2}+o(1)$.
In parallel lines, much recent progress has been made for random matrices with an inhomogeneous variance profile, in the dense case where the matrix typically has $\Theta(n^{2})$ nonzero entries. See [rudelson2016singular] for least singular value estimates in the Gaussian case and see [MR3878135], [MR3857860], [cook2article] for the general case where the circular law was proven when the matrix has a doubly stochastic variance profile. However, the case of sparse structured matrices was not well understood, and proving the convergence of its empirical spectral measure was listed as an open problem in the survey [tikhomirov2022quantitative].
For sparse structured random matrices, there are essentially only two recent papers [jain2021circular] and [tikhomirov2023pseudospectrum] studying the invertibility and pseudospectrum. The paper [jain2021circular] by Jain, Jana, Luh and O’Rourke considered a very special type of matrix model called block band matrices (see Definition 1.1), and proved convergence of empirical eigenvalue density for bandwidth $d_{n}\gg n^{33/34}$ utilizing an estimate they developed for the least singular value. Notably, the work [jain2021circular] covers discrete entry laws such as the uniform distribution on {±$\{\pm 1\}$. However, the special structure of block band matrices is utilized in an essential way, resisting a generalization to more general models such as the periodic band matrices in Definition 1.3. Another recent paper by Tikhomirov [tikhomirov2023pseudospectrum] obtained minimal singular value estimates for (the shifted version of) a wide class of random matrices with general variance profile, but assuming the random variables have a bounded density. Circular law for periodic band matrices with Gaussian entry distribution was proven as a corollary, for bandwidth up to $d_{n}\gg n^{33/34}$. The range of $d_{n}$ for circular law to hold was improved in a recent preprint [han2024circular]. However, [tikhomirov2023pseudospectrum] completely left open the case where the nonzero entries are discrete valued, say over {±$\{\pm 1\}$. And to our best knowledge, even asymptotic nonsingularity for general random band matrices with discrete entry distribution is completely unexplored. (The spectral outliers for random band matrices are much better studied; see [bandeira2024matrix] and [han2024outliers]).
The objective of this paper is to take a step into the invertibility problem for general non-Hermitian random band matrices with discrete entry distribution. We will study a class of square random matrices $A_{n}=(a_{ij})_{1\leq i,j\leq n}$ with independent entries, and that nonzero random variables are attached along a band close to the main diagonal, with the bandwidth $d_{n}$ sublinearly growing in $n$. We can also attach arbitrary integer-valued random variables at locations away from this band. Our model covers block band and periodic band matrices, although being slightly less general than the setting of [tikhomirov2023pseudospectrum] where they consider any doubly stochastic variance profile. Random band matrices, at least for the Hermitian version, are good models interpolating the de-localized phase of GOE/GUE ensembles with mean-field interaction, and the localized phase where Anderson localization takes place. The threshold for localization-delocalization transition is conjectured to be $d_{n}\sim\sqrt{n}$. A very recent breakthrough [yau2025delocalization] confirmed that when $d_{n}\gg n^{1/2+c}$, the Hermitian random band matrix exhibits complete eigenvalue delocalization, universality of local eigenvalue statistics and quantum unique ergodicity, among other things. The localized phase where $d_{n}\ll\sqrt{n}$ is also recently studied in [cipolloni2024dynamical], [chen2022random], [goldstein2022fluctuations]. A more comprehensive survey on progress in (Hermitian) random band matrices can be found in [MR3966510] and [MR3966868].
Although none of these results on Hermitian band matrices are proven for the non-Hermitian model, we can make the following postulate. The fact that a square matrix with discrete {±$\{\pm 1\}$ entry is invertible with very high probability stems from the fact that we are in the delocalized phase with strong mean-field interaction. This can be contrasted with random Schrödinger operators in the localized phase (see for example [kirsch2007invitation]), where the Wegner estimate, a quantitative version of invertibility, can generally be established only when the random potentials have a bounded density. Now we turn back to the discussion of non-Hermitian band matrices. When the band matrix entries have a bounded density, it is possible (and not hard) to prove that the matrix is almost surely invertible for bandwidth $d_{n}$ much smaller than $n^{1/2}$, but it becomes highly questionable whether the invertibility persists when the entries are discrete over {±$\{\pm 1\}$. To restate the observation, we can use the fact that a non-Hermitian random band matrix with discrete entries is invertible with probability $1-\exp(-\Omega(n^{c}))$ for some $c>0$ to convince ourselves that the band matrix still shares similar feature as a mean field model in the delocalized regime. Although we will not establish this implication in any rigorous way, we hope that this statement would provide enough intuition in our study on the invertibility of such band matrices. In particular, the main result of this paper is stated for $d_{n}\gg n^{1/2}$ (we actually need $d_{n}\gg n^{2/3}$), and we leave open the possibility of invertibility for $d_{n}\ll n^{1/2}$.
The main result of this paper is as follows.
Theorem 1.1.
Fix $\alpha\in(0,1)$ and let $d_{n}\leq n$ be a sequence of integers such that $d_{n}\geq n^{\alpha}$ when $n$ is sufficiently large.
Let $A_{n}=(a_{ij})_{1\leq i,j\leq n}$ be a family of square random matrices with independent entries. Assume that
  1. (1)
    When $|i-j|\leq d_{n}$, then each $a_{ij}$ is uniformly distributed on $\{-1,1\}$.
  2. (2)
    All the other random variables $a_{ij}$ are either identically zero or are arbitrary integer-valued random variables.
Then whenever $\alpha>\frac{2}{3}$, we have
$\mathbb{P}(A_{n}\text{ is singular})\leq\exp(-Cn^{\alpha/2}),$
for some universal constant $C>0$.
The assumption on $A_{n}$ in Theorem 1.1 is sufficiently general and covers the following two special cases as mentioned in the introduction.
Example 1.2.
(Block band matrix) Let $A_{n}$ be an $n\times n$ matrix. We say $A_{n}$ is sampled from a block band matrix ensemble if it has the form
$A_{n}=\begin{pmatrix}D_{1}&U_{2}&&&T_{m}\\ T_{1}&D_{2}&U_{3}&&\\ &T_{2}&D_{3}&\ddots&\\ &&\ddots&\ddots&U_{m}\\ U_{1}&&&T_{m-1}&D_{m}\end{pmatrix}$ (1.1)
where the unfilled sites are set zero. The blocks $D_{1},U_{1},T_{1},\cdots,D_{m},U_{m},T_{m}$ are independent copies of square random matrices with size $d_{n}$ and having i.i.d. uniform $\{-1,1\}$ entries.
Example 1.3.
(Periodic band matrix) Let $A_{n}=\{a_{ij}\}$ be an $n\times n$ square matrix. We say $A_{n}$ is sampled from a periodic band matrix model with bandwidth $d_{n}$ if $a_{ij}=0$ for any $d_{n}<|i-j|<n-d_{n}$, and $a_{ij}$ has uniform $\{-1,1\}$ distribution if $|i-j|\leq d_{n}$ or $|i-j|\geq n-d_{n}$.
We can also consider the (modified) band matrix model where $a_{ij}=0$ when $|i-j|\geq d_{n}$ and $a_{ij}$ have uniform {±$\{\pm 1\}$ distributions when $|i-j|<d_{n}$.
Remark 1.4.
The assumption in Theorem 1.1 that $a_{ij}$ are i.i.d. copies of uniform {±$\{\pm 1\}$ variables for $|i-j|\leq d_{n}$ is not significant: via a standard modification we can prove the theorem for the general case where the variables $a_{ij}$ for $|i-j|\leq d_{n}$ are independent (but not necessarily identically distributed) random variables satisfying that, for some $\beta>0$ and for any sufficiently large prime number $p\geq 2$,
$\sup_{i,j:|i-j|\leq d_{n}}\sup_{x\in\mathbb{Z}}\mathbb{P}(a_{ij}\equiv x\text{mod }p)\leq 1-\beta.$
We do not provide the details of this generalization to focus on the main ideas of proof.
Theorem 1.1 is a vast extension of the Kahn, Komlós and Szemerédi [kahn1995probability] result on invertibility of a square matrix $M_{n}$ with i.i.d. uniform {±$\{\pm 1\}$ entries, by showing that invertibility of a square random matrix still holds when we only maintain an $n^{\alpha-1}$ fraction of the total number of random variables in $M_{n}$. Before this paper, even the qualitative statement of Theorem 1.1, that $A_{n}$ is nonsingular with probability $1-o(1)$, was not proven for any band matrix model with discrete entries and sublinear bandwidth, with an exception of [jain2021circular]. In [jain2021circular], Theorem 2.1 they proved that the block band matrix model (1.1) with bandwidth $d_{n}$ is nonsingular with probability $1-C(d_{n})^{-1/2}$. Theorem 1.1 improves this singularity probability to exponential type in $n$ whenever $\alpha>\frac{2}{3}$, and yields the first proof of invertibility for periodic band matrices with {±$\{\pm 1\}$ entries for any $\alpha\in(\frac{2}{3},1)$.
We believe that proving the invertibility of general band matrices, beyond the special block band matrix setting, requires significant new ideas. The block band matrix model (1.1) in [jain2021circular] is special in that by the result of [rudelson2008littlewood], we can assume with high probability that the off-diagonal blocks $T_{1},\cdots,T_{m-1}$ and $U_{2},\cdots,U_{m}$ in (1.1) have a bounded inverse, so that we can solve each component of a vector $v$ in the kernel of $A_{n}$ using iteratively the inverse of these off-diagonal blocks. This special feature is not shared by any other model, such as the periodic band matrix model, whose invertibility is not proven before this work.
The proof of Theorem 1.1 relies on reducing the invertibility problem to a finite field, and using the inverse Littlewood-Offord counting theorems in the finite field setting. Similar ideas have appeared before in [campos2022singularity], [ferber2021counting], [campos2021singularity], but the new feature here is we have to apply the inverse counting theorem for $n/d_{n}$ times altogether to each component of a vector $v$, and the anti-concentration property for each component of the vector $v$ can be considerably different from one another.
A closely related problem to Theorem 1.1 is to estimate the least singular value of $A_{n}$, and of $zI_{n}-A_{n}$ for $z\in\mathbb{C}$ and $I_{n}$ being the identity matrix. For this purpose, an approach invented by Rudelson and Vershynin [rudelson2008littlewood] is to use the inverse Littlewood-Offord theorem on real numbers instead of the finite field, and use the notion of essential least common denominator (LCD) of a vector invented in [rudelson2008littlewood]. A major problem arises when we apply this approach to band matrices: consider for example the block band matrix model (1.1) (for a general variance profile we actually condition on some other random variables and work only with random variables in this block band structure), consider the components of a unit vector $v\in\mathbb{S}^{n-1}$ with respect to the columns indexed by $D_{1},\cdots,D_{m}$ respectively and denote each individual component by $v_{1},\cdots,v_{m}$. Then the essential LCD of each $v_{1},\cdots,v_{m}$ may vary significantly from one another in the range $[1,\exp(d_{n})]$, and the method of [rudelson2008littlewood] requires that we consider an $\epsilon_{i}$-net for each vector $v_{i}$ with $\epsilon_{i}$ depending on the essential LCD of $v_{i}$. This leads to fundamental problems since, on a row of $A_{n}$, we are using three different scales of nets to approximate, and thus only the coarsest approximation scale will be used. This leads to significant overcounting when $\epsilon_{i-1},\epsilon_{i},\epsilon_{i+1}$ are far apart from each other. Rather, working on a finite field $\mathbb{Z}_{p}$ makes the counting problem much simpler without any need of an approximation, and the finite field version of the inverse Littlewood-Offord counting theorem is far more effective in the case where $p$ is very large but the anti-concentration property of $v_{i}$ is very weak. This is the key observation that enables our argument to work through. The price we pay for this finite field approach is that we do not get an estimate on the least singular value of $A_{n}$, which is left for future study.
Finally, we conjecture that an exponential type estimate on the singularity probability of $A_{n}$ can be proven whenever $\alpha>\frac{1}{2}$ (widely perceived as the whole delocalization regime) provided that the inverse Littlewood-Offord counting problem can be significantly refined.
The invertibility of other sparse random matrix models have been studied much earlier: the sparse random graph model was considered by Rudelson and Basak in [basak2017invertibility] and [basak2021sharp], and a combinatorial model with sparsity $p$ was considered in Ferber, Kwan and Sauermann [ferber2022singularity]. For these models they prove that whenever the sparsity $p$ satisfies $p\geq(1+\epsilon)\frac{\log n}{n}$ then the matrix is invertible with probability $1-o(1)$, which is essentially best possible. However, the band matrix model considered in this paper is believed to be very different from these sparse matrix models, and the techniques do not apply here.
2. The proof
We first introduce the main technical tool for the proof of Theorem 1.1 in Section 2.1, and complete the proof of Theorem 1.1 in Section 2.2.
2.1. Finite field inverse Littlewood-Offord counting problem
We shall use one main result from [campos2022singularity] on an inverse Littlewood-Offord counting problem on finite fields. This approach of studying the inverse Littlewood-Offord counting problem on finite fields was first initiated by [ferber2021counting] and an earlier version of the result from [campos2022singularity] can be found in [campos2021singularity].
Let $p$ be an odd prime number and $\mathbb{Z}_{p}:=\mathbb{Z}/p\mathbb{Z}$ the finite field with torsion $p$. Fix an integer $d\in\mathbb{N}$. For a vector $v\in\mathbb{Z}_{p}^{d}$ and $\mu\in[0,1]$ define the random variable $X_{\mu}(v)=\epsilon_{1}v_{1}+\cdots+\epsilon_{d}v_{d}$, where $\epsilon_{i}\in\{-1,0,1\}$ are i.i.d. random variables with $\mathbb{P}(\epsilon_{i}=1)=\mathbb{P}(\epsilon_{i}=-1)=\mu/2$.
We then denote by $\rho_{\mu}(v)=\max_{x\in\mathbb{Z}_{p}}\mathbb{P}(X_{\mu}(v)=x)$, and denote by $\rho_{1}(v)=\rho(v)$. Let $|v|$ denote the number of non-zero coordinates of $v$, and for a subset $T\subset[n]$ we denote by $v_{T}:=(v_{i})_{i\in T}$ the restriction of $v$ to $T$.
For $d^{\prime}\in\mathbb{N}_{+}$ and a vector $w=(w_{1},\cdots,w_{d^{\prime}})$ we define the neighborhood of $w$ (relative to $\mu$) by
$N_{\mu}(w):=\{x\in\mathbb{Z}_{p}:\mathbb{P}(X_{\mu}(w)=x)\geq\frac{1}{2}\mathbb{P}(X_{\mu}(w)=0)\}.$
This definition of neighborhood is motivated by the fact that, when $\mu\in(0,\frac{1}{2}]$, the random walk $X_{\mu}(w)$ has the highest probability of taking the value 0 among all the values in $\mathbb{Z}_{p}$ (see [tao2006additive], Corollary 7.12). Then $N_{\mu}(w)$ collects the values $x\in\mathbb{Z}_{p}$ to which the random walk $X_{\mu}(w)$ is at least half as likely to reach compared to the value 0.
The main inverse theorem that we shall use is Theorem 2.1, taken from [campos2022singularity], Theorem 2. The original version of Theorem 2 in [campos2022singularity] actually allows for an additional parameter $k\in\mathbb{N}$. Although this parameter $k$ is fundamental for the main result in [campos2022singularity], it only leads to negligible improvements in our setting, so we only state the theorem for $k=1$.
Theorem 2.1.
Let $\mu\in(0,\frac{1}{4}]$, $d\in\mathbb{N}$, $p$ a prime number and $v\in\mathbb{Z}_{p}^{d}$. Write $D=\frac{2}{\mu}\log\rho_{\mu}(v)^{-1}$. Suppose that $|v|\geq D$ and $\rho_{\mu}(v)\geq\frac{2}{p}$. Then we can find $T\subset[d]$ with $|T|\leq D$ such that, if we denote $w=v_{T}$, then $v_{i}\in N_{\mu}(w)$ for all but at most $D$ values of $i\in[n]$. Moreover, we have
$|N_{\mu}(w)|\leq\frac{256}{\rho_{\mu}(v)}.$
We also need the following lemma 2.2 on the Fourier side, which can be found in the journal version of [campos2022singularity], Lemma 5. This type of lemma prevails in the literature, starting from [kahn1995probability], and see also [tao2006additive], Section 7.
Lemma 2.2.
For given $c>0$ and $\mu\in(0,1]$, we can find $\nu\in(0,1)$ and $K>0$ such that for all $d$-dimensional vector $v$ with $\rho_{\mu}(v)=\Omega(d/p)$ and $|v|\geq K$, we have
$\rho_{\mu}(v)\leq c\rho_{\nu}(v).$
2.2. The main argument for Theorem 1.1
We begin with some simple observations from linear algebra. For any fixed prime number $p$, $A_{n}$ is singular on $\mathbb{R}$ implies that $A_{n}$ is singular on the finite field $\mathbb{Z}_{p}$ since $A_{n}$ has integer coefficients. Then $A_{n}$ is singular on $\mathbb{Z}_{p}$ implies that there exists some index $\mathbf{I}\in[n]$ such that the $\mathbf{I}$-th row of $A_{n}$ lies in the $\mathbb{Z}_{p}$-linear span of all the other rows of $A_{n}$. We will make a careful choice of the prime number $p$, which roughly has order $\exp(n^{\alpha/2})$, and study the singularity probability of $A_{n}$ on $\mathbb{Z}_{p}$. The latter probability yields an upper bound for the singularity probability of $A_{n}$ on $\mathbb{R}$.
From this observation, we let $A_{n}^{\mathbf{I}}$ denote the matrix obtained from $A_{n}$ by setting its $\mathbf{I}$-th row to be zero, and denote by $v^{\mathbf{I}}$ the $\mathbf{I}$-th row of $A_{n}$. The rationale here is that, if $v^{\mathbf{I}}$ lies in the span of the rows of $A_{n}^{\mathbf{I}}$, then $v^{\mathbf{I}}$ must be orthogonal to any nonzero vector $v\in\ker(A_{n}^{\mathbf{I}})$ (where orthogonality means that $\langle v^{\mathbf{I}},v\rangle$=0 on $\mathbb{Z}_{p}$). By independence of $v^{\mathbf{I}}$ and $A_{n}^{\mathbf{I}}$, we only need to determine certain arithmetic structure for any nonzero vectors $v\in\ker A_{n}^{\mathbf{I}}\subset\mathbb{Z}_{p}^{n}$.
In contrast to the full matrix setting (a square matrix with all i.i.d. entries), the main technical difficulty for random band matrices is that the vector $v^{\mathbf{I}}$ may only have $2d_{n}\ll n$ coordinates with nonzero random variables, and thus we have to show that any nonzero vector $v\in\ker A_{n}^{\mathbf{I}}$ is nonzero when restricted to these $2d_{n}$ coordinates and the restriction of $v$ to these entries has a sufficiently good anti-concentration property. When $d_{n}=\Theta(n)$, structure theorems can be found in [rudelson2016singular] and [cook2018lower]. But when $d_{n}$ is sublinear in $n$, no structural results for $v\in\ker A_{n}^{\mathbf{I}}$ were previously available for a general variance profile.
Before stating our main structure theorem on vectors in $\ker(A_{n}^{\mathbf{I}})$, we introduce some notations on decomposition of the structure of $A_{n}$. Although we have allowed for a rather arbitrary variance profile for $A_{n}$ in Theorem 1.1, the first step of the proof is to replace $A_{n}$ by another matrix with a block band structure as in (1.1).
We define $e_{n}=\lfloor d_{n}/s\rfloor$ (for some $s\in(2,3)$ to be later fixed) and consider a partition of $[1,n]$ by intervals of length $e_{n}$:
$[1,n]=\cup_{k=1}^{\lfloor n/e_{n}\rfloor+1}I_{k},\quad I_{k}=\begin{cases}[(k-1)e_{n}+1,ke_{n}],\quad\forall 1\leq k\leq\lfloor n/e_{n}\rfloor\\ [\lfloor n/e_{n}\rfloor e_{n}+1,n],\quad k=\lfloor n/e_{n}\rfloor+1,\end{cases}$
where we choose some $s\in(2,3)$ such that $n-e_{n}\lfloor n/e_{n}\rfloor\geq e_{n}/2$. This ensures that the shortest length of the intervals among $I_{k}$ is at least half as long as the other intervals among $I_{k}$. Let $n_{\mathbf{I}}\in\mathbb{N}$ be such that $(n_{\mathbf{I}}-1)e_{n}+1\leq\mathbf{I}\leq n_{\mathbf{I}}e_{n}$, so that $n_{\mathbf{I}}$ is the subscript $k$ of the sub-interval $I_{k}$ where $\mathbf{I}\in I_{k}$.
For this subscript $\mathbf{I}$ and $n_{\mathbf{I}}$, we define the following matrix from $A_{n}=(a_{ij})$ that will be used for comparison:
$\mathcal{D}_{\mathbf{I}}$ :=$:=\{a_{ij}\mathbf{1}_{i,j\in I_{k}\text{ or }i\in I_{k-1},j\in I_{k}\text{ or }i\in I_{k},j\in I_{k-1}\text{ for some }k=1,\cdots,\lfloor n/d_{n}\rfloor+1}\}\cdot\mathbf{1}_{i\neq\mathbf{I}}.$ (2.1)
(where we interpret $I_{0}=J_{0}=\emptyset$).
In matrix form, we can write $\mathcal{D}_{\mathbf{I}}$ in the following block form:
$\mathcal{D}_{\mathbf{I}}=\begin{bmatrix}D_{1}&U_{2}&0&\cdots&0\\ T_{1}&D_{2}&U_{3}&\cdots&0\\ 0&T_{2}&D_{3}&\cdots&0\\ \ \vdots&\vdots&\vdots&\ddots&U_{\lfloor n/e_{n}\rfloor+1}\\ 0&0&0&T_{\lfloor n/e_{n}\rfloor}&D_{\lfloor n/e_{n}\rfloor+1}\end{bmatrix},$ (2.2)
where by the assumption of Theorem 1.1, the entries in the blocks $D_{i},U_{i},T_{i},1\leq i\leq\lfloor n/e_{n}\rfloor+1$ are independent and uniformly distributed over {±$\{\pm 1\}$ except the $\mathbf{I}$-th row of $\mathcal{D}_{\mathbf{I}}$ which is identically 0. In other words, $\mathcal{D}_{\mathbf{I}}$ is a matrix extracted from $A_{n}$ by zeroing out its $\mathbf{I}$-th row and zeroing out all the entries not in the main diagonal and first off-diagonal blocks labeled by $I_{k},k=1,\cdots,\lfloor n/e_{n}\rfloor+1$
The main structural theorem on $\ker A_{n}^{\mathbf{I}}$ is stated as follows:
Theorem 2.3.
In the setting of Theorem 1.1, there exist constants $\rho>0$, $\tau>0$ such that the following holds with $p=\exp(\rho n^{\alpha/2})$:  [^1] By the prime number theorem, there must exist a prime number in $[\exp(\frac{\rho}{2}n^{\alpha/2}),\exp(\rho n^{\alpha/2})]$, and we simply use $p:=\exp(\rho n^{\alpha/2})$ to denote any one of these prime numbers.
On the finite field $\mathbb{Z}_{p}$, with probability at least $1-\exp(-\Omega(n))$, the following is true for all nonzero vectors $v\in\ker A_{n}^{\mathbf{I}}\subset\mathbb{Z}_{p}^{n}$: if we denote by $v_{I_{n_{\mathbf{I}}}}$ the restriction of $v$ to the interval $I_{n_{\mathbf{I}}}$, then we have, with $\rho(v)$ defined in Section 2.1,
$\rho(v_{I_{n_{\mathbf{I}}}})\leq\exp(-\tau n^{\alpha/2}).$
The proof of Theorem 2.3 in fact shows that $\rho(v_{I_{k}})\leq\exp(-\tau n^{\alpha/2})$ for each $k\leq\lfloor n/e_{n}\rfloor+1$, that is, we obtain strong arithmetic properties for each component of $v$ on an interval with sublinear length.
The proof of Theorem 1.1 is immediate from Theorem 2.3:
Proof of Theorem 1.1.
By the discussion at the beginning of Section 2.2, we have
$\mathbb{P}(\det A_{n}=0)\leq\mathbb{P}(\det A_{n}=0\text{ on }\mathbb{Z}_{p})\leq\sum_{\mathbf{I}=1}^{n}\mathbb{P}(v^{\mathbf{I}}\in\operatorname{Span}((v^{j})_{j\neq\mathbf{I}})\text{ on }\mathbb{Z}_{p})$
$\leq\sum_{\mathbf{I}=1}^{n}\mathbb{P}(\langle v^{\mathbf{I}},v\rangle=0\text{ for some }v\neq 0:v\in\ker A_{n}^{\mathbf{I}}\subset\mathbb{Z}_{p}^{n})$
$\leq\sum_{\mathbf{I}=1}^{n}\min_{0\neq v\in\ker A_{n}^{\mathbf{I}}}\rho(v_{I_{n_{\mathbf{I}}}})\leq n\exp(-\Omega(n^{\alpha/2}))+n\exp(-\Omega(n)),$
where the last inequality follows from conditioning on the random variables of $v^{\mathbf{I}}$ indexed by $[n]\setminus I_{n_{\mathbf{I}}}$ and partitioning the event into two cases: either the conclusion of Theorem 2.3 holds for all $v_{I_{n_{\mathbf{I}}}}$ and all $\mathbf{I}$, or the conclusion of the theorem does not hold for one choice of $\mathbf{I}$, and the latter probability is at most $n\exp(-\Omega(n))$. This concludes the proof. ∎
Then we begin with the proof of Theorem 2.3.
Proof of Theorem 2.3.
Let $\mu\in(0,\frac{1}{4}]$ be a constant to be determined later, and $v\in\mathbb{Z}_{p}^{n}\setminus\{0\}$. We consider a dyadic decomposition of the range $\rho_{\mu}(v_{I_{1}}),\rho_{\mu}(v_{I_{2}}),\cdots,\rho_{\mu}(v_{I_{\lfloor n/e_{n}\rfloor+1}})\in(0,\frac{2}{p}]\cup[\frac{2}{p},1]$ via the following rule: for each $k,1\leq k\leq\lfloor n/e_{n}\rfloor+1$,
  1. (1)
    If $0<|v_{I_{k}}|\leq K$ where $K$ is given in Lemma 2.2, then we only need at most $K$ coordinates to define $v_{I_{k}}$ and no dyadic decomposition for the range of $\rho(v_{I_{k}})$ is needed.
  2. (2)
    If $|v_{I_{k}}|\geq K$ and $\rho_{\mu}(v_{I_{k}})\leq\frac{2}{p}$, then again we do not take a dyadic decomposition for the range of $\rho_{\mu}(v_{I_{k}})$ and $v_{I_{k}}$ can take all the $p^{|I_{k}|}$ possible vectors.
  3. (3)
    In the remaining case where $|v_{I_{k}}|\geq K$ and $\rho_{\mu}(v_{I_{k}})\in[\frac{2}{p},1]$, we consider a dyadic decomposition for the interval $[\frac{2}{p},1]\subset\cup_{i=1}^{\log_{2}p}[2^{-i},2^{-i+1}]$ for the range of $\rho_{\mu}(v_{I_{k}})$.
The total number of intervals needed for the dyadic decomposition is $O(\log_{2}p)^{n/e_{n}}=2^{o(n)}$ under the assumption that $e_{n}\geq n^{\alpha}$ and $\alpha>\frac{2}{3}$. ∎
The proof of this theorem follows from exhausting all the possible values for $\rho_{\mu}(v_{I_{k}}),1\leq k\leq\lfloor n/e_{n}\rfloor+1$ and using the inverse Littlewood-Offord counting theorem, Theorem 2.1, for vectors in case (3). We first consider two special cases:
(1) Assume that $\rho_{\mu}(v_{I_{k}})\leq\frac{2}{p}$ for all $1\leq k\leq\lfloor n/e_{n}\rfloor+1:k\neq n_{\mathbf{I}}$. That is, all components $v_{I_{k}}$ of $v$, except the $k=n_{\mathbf{I}}$-th component that we ask about, have very good anti-concentration properties. Then we can upper bound $\mathbb{P}(A_{n}^{\mathbf{I}}v=0)$ as follows: for each row of $A_{n}^{\mathbf{I}}$ whose index lies in $I_{k}$, $k\neq n_{\mathbf{I}}$, we use the randomness in the $D_{k}$ component in (2.2) and condition on all other random variables in the same row. For the rows with index lying in $I_{n_{\mathbf{I}}}$, we either use the randomness in $U_{k+1}$ or the randomness in $T_{k-1}$ 222we can choose either of $U_{k+1}$ or $T_{k-1}$ if $k\neq 1$ and $k\neq\lfloor n/e_{n}\rfloor+1$, but we can only choose $U_{2}$ for $k=1$ and we can only choose $T_{\lfloor n/e_{n}\rfloor}$ for $k=\lfloor n/e_{n}\rfloor+1$. Then
$\mathbb{P}(A_{n}^{\mathbf{I}}v=0)\leq\prod_{k\leq\lfloor n/e_{n}\rfloor+1,k\neq n_{\mathbf{I}}}\rho(v_{I_{k}})^{|I_{k}|}\cdot\max_{k\in\{n_{\mathbf{I}}-1,n_{\mathbf{I}}+1\}}\rho(v_{I_{k}})^{|I_{n_{\mathbf{I}}}|-1}\leq(\frac{2}{p})^{n-1},$
where we use $\rho(v_{I_{k}})\leq\rho_{\mu}(v_{I_{k}})$. Let $\tau>0$ be another constant to be fixed later, and we upper bound the cardinality of all vectors $v_{I_{n_{\mathbf{I}}}}\in\mathbb{Z}_{p}^{|n_{\mathbf{I}}|}$ with $\rho_{\mu}(v_{I_{n_{\mathbf{I}}}})\geq\exp(-\tau n^{\alpha/2})$ by
$2^{|I_{n_{\mathbf{I}}}|}p^{\frac{2\tau}{\mu}n^{\alpha/2}}\cdot p^{\frac{2\tau}{\mu}n^{\alpha/2}}\cdot(256\exp(\tau n^{\alpha/2}))^{|I_{n_{\mathbf{I}}}|}$
where we apply Theorem 2.1. The $2^{|I_{n_{\mathbf{I}}}|}$ factor upper bounds the choice for $T$, the first $p^{\frac{2\tau}{\mu}n^{\alpha/2}}$ term upper bounds the number of vectors $v_{T}$, the second such term upper bounds the number of choice of $v$ in at most $D:=\frac{2\tau}{\mu}n^{\alpha/2}$ coordinates not included in $N_{\mu}(w)$ with $w:=v_{T}$, and the last term bounds the number of choices of coordinates of $v_{i}\in N_{\mu}(w)$. (The case where $|v|\leq\max(D,K)$, which is excluded by Theorem 2.1, is also upper bounded by the above expression.)
Then combining the previous two expressions and applying a union bound, we get
( $\mathbb{P}(\text{ There exists }v\in\ker A_{n}^{\mathbf{I}}:\rho_{\mu}(v_{I_{k}})\leq\frac{2}{p}\forall k\neq n_{\mathbf{I}},\quad\rho_{\mu}(v_{I_{n_{\mathbf{I}}}})\geq\exp(-\tau n^{\alpha/2}))$ (2.3)
$\leq 2^{n}p^{-n+1}\cdot p^{n-|I_{n_{\mathbf{I}}}|}\cdot 2^{O_{\mu,\tau}(n^{\alpha/2}\log n)}\exp(\tau n^{\alpha/2}|I_{n_{\mathbf{I}}}|).$
where $a=O_{\mu,\tau}(b)$ means $a\leq Cb$ for $C$ a constant depending only on $\mu,\tau$ and $\alpha$.
Recall that we take $p=\exp(\rho n^{\alpha/2})$ and $|I_{n_{\mathbf{I}}}|\geq n^{\alpha}$ with $\alpha>\frac{2}{3}$, we see that as long as $\tau<\rho/2$, the above expression is super-exponentially small, which is actually at most $\exp(-\Omega(n^{1.5\alpha}))$. Thus with probability at least $1-\exp(-\Omega(n^{1.5\alpha}))$, no such vector as described in the first line of (2.3) lies in $\ker A_{n}^{\mathbf{I}}$.
(2) Then we consider the second case where none of the $v_{I_{k}}$, $11\leq k\leq\lfloor n/e_{n}\rfloor+$ 1 satisfies $\rho_{\mu}(v_{I_{k}})\leq\frac{e_{n}}{p}$. We first assume that $|v_{I_{k}}|\geq K$ for each $K$ (and treat the other case $|v_{I_{k}}|\leq K$ at the end of step (2) by a simple modification). Let $t_{1},\cdots,t_{\lfloor n/e_{n}\rfloor+1}$ be integers such that $\rho_{\mu}(v_{I_{k}})\in[2^{-t_{k}},2^{-t_{k}+1}]$ for each $k$. Then we upper bound the probability $\mathbb{P}(A_{n}^{\mathbf{I}}v=0)$ by conditioning on all the random variables away from the diagonal blocks $D_{1},D_{2},\cdots,D_{\lfloor n/e_{n}\rfloor+1}$ and using randomness only from these diagonal blocks:
$\mathbb{P}(A_{n}^{\mathbf{I}}v=0)\leq\prod_{k=1}^{\lfloor n/e_{n}\rfloor+1}|\rho(v_{I_{k}})|^{|I_{k}|-1_{k=n_{\mathbf{I}}}}$
$\leq c^{n-1}\prod_{k=1}^{\lfloor n/e_{n}\rfloor+1}|\rho_{\mu}(v_{I_{k}})|^{|I_{k}|-1_{k=n_{\mathbf{I}}}}\leq\frac{p}{c}\cdot c^{n}\prod_{k=1}^{\lfloor n/e_{n}\rfloor+1}(2^{-t_{k}+1})^{|I_{k}|},$
where in the second line we take any value of $c\in(0,1)$ to be fixed at the end of proof and use Lemma 2.2 to get a value $\mu\in(0,\frac{1}{4}]$ such that Lemma 2.2 holds with this value $c$. The indicator function $1_{k=n_{\mathbf{I}}}$ comes from removal of the $\mathbf{I}$-th row from $A_{n}^{\mathbf{I}}$.
Then we compute the cardinality of vectors $v_{I_{1}},\cdots,v_{I_{k}},\cdots,v_{I_{\lfloor n/e_{n}\rfloor+1}}$ satisfying the stated constraint on anticoncentration. Set $D_{k}=\frac{2}{\mu}(t_{k}+1)$ for each $k$, as we apply Theorem 2.1, for each $v_{I_{k}}$ we have at most $2^{|I_{k}|}$ choices for $T$, $p^{D_{k}}$ choices for $w$, and $2^{(t_{k}+9)}$ choices for every other coordinate of $v_{I_{k}}$. The total cardinality for these vectors of $v_{I_{k}}$ is at most
$2^{|I_{k}|+9}p^{D_{k}}(2^{\mu D_{k}/2})^{|I_{k}|-D_{k}}\leq 2^{|I_{k}|+9}\cdot 2^{(t_{k}+1)|I_{k}|}\cdot e^{\frac{2}{\mu}(\log p)^{2}},$
where the term $e^{\frac{2}{\mu}(\log p)^{2}}$ comes from optimizing $(\frac{p}{2^{\mu D_{k}/2}})^{D_{k}}\leq(\frac{p}{2^{t_{k}+1}})^{\frac{2}{\mu}(t_{k}+1)}$ over $t_{k}\geq 1$.
Taking the product for all $k=1,2,\cdots,\lfloor n/e_{n}\rfloor+1$, we have via a direct union bound,
$\mathbb{P}(\text{There exists }v\in\ker A_{n}^{\mathbf{I}}:\rho_{\mu}(v_{I_{k}})\in[2^{-t_{k},2^{-t_{k}+1}}]\forall k,|v_{I_{k}}|\geq K)$ (2.4)
$\leq pc^{n-1}2^{20n}e^{\frac{2}{\mu}(\log p)^{2}\frac{n}{e_{n}}}.$
By our assumption, $(\log p)^{2}\frac{n}{e_{n}}\leq\rho^{2}n^{\alpha}\frac{n}{n^{\alpha}/s}\leq\rho^{2}sn$ for $s$ fixed before the statement of Theorem 2.3, so that by setting $c>0$ sufficiently small (this will increase the value of $\mu$) and then setting $\rho>0$ sufficiently small relative to $\mu$, we can guarantee that the second lie of (2.4) is at most $\exp(-\Omega(n))$. Since the number of dyadic decompositions is at most $2^{o(n)}$ by the beginning of the proof, the eventual probability for a vector in case (2) to be in $\ker A_{n}^{\mathbf{I}}$ is $\exp(-\Omega(n))$.
We now treat the leftover case where for some $k$, we have $v_{I_{k}}=0$ or $0<|v_{I_{k}}|<K$. For the first case we only need trivial modification in the above computation, and for the second case we use the trivial fact that $\rho(v_{I_{k}})\leq\frac{1}{2}$ combined with the upper bound on the number of such $v_{I_{k}}$: the number of these vectors is at most $(e_{n})^{K}p^{K}=e^{O(n^{\alpha/2}\log n)}$. Then the contribution of these vectors to the union bound of (2.4) is at most $2^{-e_{n}}e^{O(n^{\alpha/2}\log n)}=2^{-\Omega(e_{n})}$, and a similar estimate as in (2.4) can be proven to be at most $\exp(-\Omega(n))$.
(2’) We also need to consider the case where for exactly one $k$ we have $\rho_{\mu}(v_{I_{k}})\leq\frac{e_{n}}{p}$ holds. For this case we use the same estimate as in case (2) for $v_{I_{l}},l\neq k$, but for component $v_{I_{k}}$ we use all the $p^{|I_{k}|}$ possible values of $v_{I_{k}}$. Then the probability estimate in (2.4) would now have the form $pc^{n-1-|I_{k}|}2^{20n}e^{\frac{2}{\mu}(\log p)^{2}\frac{n}{e_{n}}}\cdot(e_{n})^{e_{n}}$. But we have $(e_{n})^{e_{n}}=2^{o(n)}$ (we shall assume $e_{n}$ is truly power-law in $n$ with $\alpha<1$), so the above quantity is again $\exp(-\Omega(n))$.
(3) Finally, we consider the general case where $\rho_{\mu}(v_{I_{n_{\mathbf{I}}}})\geq\exp(-\tau n^{\alpha/2})$, with $\tau=\rho/2$, and there exists some $k\neq n_{\mathbf{I}},k\neq\lfloor n/e_{n}\rfloor+1$ such that $\rho_{\mu}(v_{I_{k}})\leq\frac{e_{n}}{p},\quad|v_{I_{k}}|\geq K$. Observe that the complement of case (3) is studied in case (2) and (2’). Fix one such value $k_{*}$ such that $\rho_{\mu}(v_{I_{k_{*}}})\leq\frac{e_{n}}{p},\quad|v_{I_{k_{*}}}|\geq K$. We assume without loss of generality that $k_{*}<n_{\mathbf{I}}$.
Then we can find a sequence of integers $k_{1}>k_{2}\geq\cdots>k_{*}$ such that $k_{i}$ is the largest index in $[1,k_{i-1}]$ such that $\rho_{\mu}(v_{I_{k_{i}}})\leq\rho_{\mu}(v_{I_{k_{i-1}}})$ for each $i$, where we denote by $k_{0}=n_{\mathbf{I}}$. That is, we find the maximal chain of indices $k_{0},k_{1},\cdots,k_{*}$ where $\rho_{\mu}(v_{I_{k_{i}}})$ is increasing in $i$. We clearly have by definition of $n_{\mathbf{I}}$ and $k_{*}$ that
$\exp(-\rho n^{\alpha/2}/4)\geq\frac{\rho_{\mu}(v_{I_{k_{*}}})}{\rho_{\mu}(v_{I_{n_{\mathbf{I}}}})}=\prod_{i=1}^{*}\frac{\rho_{\mu}(v_{I_{k_{i}}})}{\rho_{\mu}(v_{I_{k_{i-1}}})}\geq\prod_{i=1}^{*}\frac{\rho_{\mu}(v_{I_{k_{i}}})}{\rho_{\mu}(v_{I_{k_{i}+1}})}.$ (2.5)
We upper bound $\mathbb{P}(A_{n}^{\mathbf{I}}v=0)$ via the following approach: for rows indexed in $I_{k_{i}+1}$ we use randomness from $T_{I_{k_{i}}}$ (see (2.2)), and for all other rows indexed by $I_{l}$ we use randomness from the diagonal component $D_{l}$. This helps us to make the best use of anticoncentration properties for each component of $v$.
We proceed as in case (2), take a dyadic decomposition of the range of $\rho_{\mu}(v_{I_{k}})\subset[2^{-t_{k}},2^{-t_{k}+1}]$ whenever $\rho_{\mu}(v_{I_{k}})\geq\frac{e_{n}}{p}$, and simply take the upper bound $\rho_{\mu}(v_{I_{k}})\leq\frac{e_{n}}{p}$ in the other case. The cardinality of each vector component $v_{I_{k}}$ is again bounded via applying Theorem 2.1. Then we have, using the simple estimate $\rho(v_{I_{k}})\leq\rho_{\mu}(v_{I_{k}})$ for any $k$ and $\mu\in(0,\frac{1}{4}]$,
$\mathbb{P}(A_{n}^{\mathbf{I}}v=0)\leq\prod_{i=1}^{\lfloor n/e_{n}\rfloor+1}|\rho_{\mu}(v_{I_{i}})|^{|I_{i}|}\cdot\prod_{i=1}^{*}\left|\frac{\rho_{\mu}(v_{I_{k_{i}}})}{\rho_{\mu}(v_{I_{k_{i}+1}})}\right|^{|I_{k_{i}+1}|}.$ (2.6)
Then by a direct union bound, we have for this fixed dyadic partition (fixed $t_{1},\cdots,t_{k},\cdots$),
(There exists $\mathbb{P}(\text{There exists }v\in\ker A_{n}^{\mathbf{I}}\text{ with the specified property})$
$\leq 2^{20n}e^{\frac{2}{\mu}(\log p)^{2}\frac{n}{e_{n}}}\prod_{k:\rho_{\mu}(v_{I_{k}})\geq\frac{2}{p}}2^{(t_{k}+1)|I_{k}|}\prod_{k:\rho_{\mu}(v_{I_{k}})\leq\frac{2}{p}}p^{|I_{k}|}$
$\quad\cdot\prod_{k=1}^{\lfloor n/e_{n}\rfloor+1}|\rho_{\mu}(v_{I_{k}})|^{|I_{k}|}\cdot\exp(-\rho n^{\alpha/2}|I_{n_{\mathbf{I}}}|/4),$
where we have applied (2.5) to the last term on the right hand side of (2.6).
Simplifying the above expression and noting that $n^{\alpha/2}|I_{n_{\mathbf{I}}}|\geq n^{1.5\alpha}$ and $\alpha>\frac{2}{3}$, the last display is at most $\exp(-\Omega(n^{1.5\alpha}))$ and hence negligible after taking the union bound over the dyadic decomposition.
This completes the entire proof of Theorem 2.3.
2.3. Concluding remarks
Carefully checking the computations in case (2) of the proof of Theorem 2.3, we see that $\alpha>\frac{2}{3}$ is the threshold value for the whole argument to work. It might be possible to improve the restriction on $d_{n}$ by a $\sqrt{\log n}$ factor, but getting towards $\alpha<\frac{2}{3}$ may require complete new ideas. Indeed, the inverse Littlewood-Offord theorem on finite fields is well known (see [campos2021singularity], [campos2022singularity], [ferber2021counting]) to work well up to $p\leq\exp(\sqrt{n\log n})$ where $n$ is the dimension, which leads to the constraint $\alpha>\frac{2}{3}$ in our band matrix case.
Ideally, we expect our method to work well for all $\alpha\in(\frac{1}{2},1)$ (which coincides with the conjectured parameter range of delocalization in Anderson localization theory), and the range $\alpha\in(\frac{1}{2},\frac{2}{3})$ is only reachable when we can prove a much more refined version of inverse Littlewood-Offord inequality than the one presented here.
There have been many papers studying integral random matrices over a finite field, see for example [fulman2015stein], [nguyen2022random] and references therein. This paper appears to be the first to consider the rank of a truly structured integral band matrix over a finite field (but see the recent work of Mészáros [meszaros2024phase] on a random band matrix whose nonzero entries are uniform generated over the $p$-adic integers, and the behavior of the model in [meszaros2024phase] is very different from the model in this paper). In all previous papers considering a square (symmetric) or rectangular matrix on a finite field, the law of the rank of these random matrices often has a universal limiting distribution (see [ferber2023random] [eberhard2022characteristic]) for all prime $p$ growing not too fast in $n$, and in particular we can prove asymptotic non-singularity of these matrices by taking $p\to\infty$ arbitrarily slowly. However, in our band matrix setting, we cannot gain any information on the invertibility of $A_{n}$ when $p$ does not grow fast enough in $n$, and we also cannot work through the proof when $p$ grows too quickly in $n$. Thus, one really needs a careful choice of $p$ as a function in $n$ in the band matrix setting.
Funding
The author is supported by a Simons Foundation Grant (601948, DJ).
[…]